6. Researchers took random samples of right-handed men and women and determined the numbers whose feet were the same size (L-R), the left foot larger than the right (L > R) and the number where the left foot was smaller than the right (L < R) Do the following data indicate that gender has an effect on the development of foot asymmetry? Men 2 10 28 Women 55 18 14 (a) State appropriate null and alternate hypotheses. (b) Compute the value of an appropriate test statistic (c) State your conclusions. (d) What is the approximate p-value? C1] L2] 01] x3(2) = 4.61, χる5(2) = 5.99, χ2 25(2) = 7.38, χ2 1 (2) = 9.21, χ2 05(2) = 10.59; x3 (3) = 6.25, 2 5(3) = 7.81, χ225(3) = 9.35,墻1(3) = 11.34, χ·005(3)-12.84; x3 (4) = 7.78, χ2 5(4) = 9.49, χ2 25(4) = 11.14, xh(4) = 13.27, χ205(4) = 14.86.
Answer #1

(a)

H0: Null Hypothesis: Gender has no effect on the development of foot asymmetry.

HA: Alternative Hypothesis: Gender has effect on the development of foot asymmetry.

(b)

Observed frequencies:

L>R L=R L<R Total
Men 2 10 28 40
Women 55 18 14 87
Total 57 28 42 127

Expected frequencies:

L>R L=R L<R Total
Men 57X40/127=17.95 8.82 13.23 40
Women 39.05 19.18 28.77 87
Total 57 28 42 127

The Test statistic (\chi ^{2}) is calculated as follows:

O E (O - E)2/E
2 17.95 14.18
10 8.82 0.16
28 13.23 16.50
55 39.05 6.52
18 19.18 0.07
14 28.77 7.58
Total = \chi ^{2} = 45.00

Test statistic = \chi ^{2} = 45.00

(c)

Take \alpha = 0.05

ndf = (r - 1) X (c - 1)

= (3 - 1) X (2 - 1) = 2

From Table, critical value of \chi ^{2} = 5.99

Sinc the calculated value of \chi ^{2} = 45.00 is greater than critical value of \chi ^{2} = 5.99, the difference is significant. Reject null hypothesis.

Conclusion:

The data support the claim that Gender has effect on the development of foot asymmetry.

(d)

Test statistic = \chi ^{2} = 45.00

ndf = 2

By Technology, p - value = 0.00001